OpenStudy (anonymous):

Sets and probability Suppose n(A upside down U B)=5, n(A U B)=35, and n(A)=12 What is n(B)?

OpenStudy (watchmath):

\(n(A\cup B)=n(A)+n(B)-n(A\cap B)\) \(35=12+n(B)-5\) \(n(B)=28\).